Future value of ordinary annuity


what is the future value of an annuity if the size of each payment is 20,000 payable at the end of each year for 3years at an interest of 9% compounded Annually ?


1
Expert's answer
2021-03-21T20:06:58-0400

"FVA_{Ordinary}=P\times\frac{[(1+r)^{n}-1]}{i}"

"FVA_{Ordinary}=20,000\times\frac{[(1+0.09)^{3}-1]}{0.09}=20,000\times3,2781=65,562"